### v5

@@ D.1 Notation
[…]
@@ D.2 Theorem T-2 in HP
$\bm{K}_{t}=\bm{I}_{S}-\bm{B}_{t}^{\dagger}\bm{B}_{t}\;\text{ and }\;{\bm{\mathcal{K}}}=\bm{I}_{S}-{\bm{\mathcal{B}}}^{\dagger}{\bm{\mathcal{B}}},$
where $\bm{I}_{S}$ denotes the identity matrix of dimension $S$. Note that ${\bm{\mathcal{K}}}$ and $\bm{K}_{t}$ are orthogonal projection matrices in $\mathbb{R}^{S}$ that only depend on the binary matrices ${\bm{\mathcal{B}}}$ and $\bm{B}_{t}$. Theorem T-2 in HP shows that
@@
(D.2) $\displaystyle\bm{Q}_{S}(t)$ $\displaystyle=\bm{B}_{t}^{\dagger}\bm{Q}_{Z}(t)+\bm{K}_{t}\tilde{\bm{\lambda}},$
@@
D.3 Identification in the $2\times 3$ model
@@
$\displaystyle{\bm{\mathcal{B}}}$ $\displaystyle=\begin{bmatrix}1&0&0&1&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&1&0&1&1\\ 0&0&1&0&1\\ 0&0&1&0&0\end{bmatrix}.$
@@
$\displaystyle{\bm{\mathcal{B}}}^{\dagger}$ $\displaystyle=\frac{1}{6}\begin{bmatrix}1&5&1&-1&1&-1\\ -1&1&5&1&-1&1\\ 1&-1&1&-1&1&5\\ 4&-4&-2&2&-2&2\\ -2&2&-2&2&4&-4\end{bmatrix},$
this is not very transparent, however. To derive our identification results, we use (D.2) instead. Note that the equation $\bm{Q}_{S}(t)=\bm{B}_{t}^{\dagger}\bm{Q}_{Z}(t)+\bm{K}_{t}\tilde{\bm{\lambda}}$ holds for any function of $Y(t)$. If we take it to be a constant function of $Y(t)$, we get $\bm{Q}_{S}(t)=\mathbb{E}\bm{S}=\bm{P}_{S}$ and $\bm{Q}_{Z}(t)=\bm{P}_{Z}(t)$, so that (D.2) boils down to
(D.3) $\bm{P}_{S}=\bm{B}_{t}^{\dagger}\bm{P}_{Z}(t)+\bm{K}_{t}\tilde{\bm{\lambda}}\;\text{ for all values of }t.$
@@
Let $(\bm{e}_{s})_{s=1,\ldots,5}$ denote the standard basis vectors in $\mathbb{R}^{5}$. If $\bm{e}^{\prime}_{s}\bm{K}_{t}=0$, then (D.3) point-identifies $\Pr(i\in C^{s})=\bm{e}^{\prime}_{s}\bm{P}_{S}=\bm{e}^{\prime}_{s}\bm{B}_{t}^{\dagger}\bm{P}_{Z}(t)$. Clearly,
@@
this reproduces our identification results for $\bm{P}_{S}$ in Proposition 5:
@@
Returning to the counterfactual outcomes $Y(t)$, the same argument results in Proposition 6:
@@
D.4 Identification in the $3\times 3$ model
@@
Let $(\bm{e}_{s})_{s=1,\ldots,8}$ denote the standard basis vectors in $\mathbb{R}^{8}$. Since $\bm{K}_{0}$ has no zero column, none of the $\bm{e}^{\prime}_{s}\bm{K}_{0}$ is zero and the argument in Section D.3 show that no population share $\Pr(i\in C^{s})$ is point-identified by $\bm{B}_{0}^{\dagger}\bm{P}_{Z}(0)$. On the other hand,
@@
and $\bm{e}^{\prime}_{3}\bm{K}_{2}=\bm{e}^{\prime}_{8}\bm{K}_{2}=0$ so that the shares of $C_{222}$ and $C_{212}$ are point-identified. On the other hand, the shares of $C_{010}$, $C_{002}$, and $C_{012}$ are not identified. The results in Propositions 7 and 8 follow.
@@
where we indicate the relevant columns and rows of matrix $\bm{B}_{0}$. Given this pattern, Theorem T-3 and Remark 6.3 in Heckman and Pinto (2018) imply that the unordered monotonicity assumption is not satisfied for the $3\times 3$ model. As mentioned in the main text, this , switching from instrument value from 1 to 2 causes observations in $C_{010}$ to move to treatment 00 0, while those in $C_{002}$ move out of treatment 00 0. Recall that the ARUM structure rules out “direct two-way flows” (that is, instrument values 11 11 and 22 22, respectively, make treatments 11 11 and 22 22 more favorable for everyone). However, the $3\times 3$ model allows for “indirect two-way flows”, where treatment 0 is not targeted by either $z=1$ or $z=2$. Unordered monotonicity is more restrictive than ARUM in that it rules out both direct and indirect two-way flows.
Appendix E The $3\times 3$ Model of Pinto (2021)
@@
(E.1) $\displaystyle\begin{split}\Pr(C_{hhh})&=P(t_{h}|z_{8}),\\ \Pr(C_{lll})&=P(t_{l}|z_{c}),\\ \Pr(C_{mmm})&=P(t_{m}|z_{e}),\\ \Pr(C_{hhm})&=P(t_{h}|z_{e})-P(t_{h}|z_{8}),\\ \Pr(C_{hll})&=P(t_{l}|z_{8})-P(t_{l}|z_{c}),\\ \Pr(C_{mlm})&=P(t_{m}|z_{c})-P(t_{m}|z_{e}),\\ \Pr(C_{hlm})&=1-P(t_{h}|z_{e})-P(t_{l}|z_{8})-P(t_{m}|z_{c}).\end{split}$
@@
(E.2) $\displaystyle\begin{split}P(t_{h}|z_{e})&\geq P(t_{h}|z_{8}),\\ P(t_{l}|z_{8})&\geq P(t_{l}|z_{c}),\\ P(t_{m}|z_{c})&\geq P(t_{m}|z_{e}),\\ 1&\geq P(t_{h}|z_{e})+P(t_{l}|z_{8})+P(t_{m}|z_{c}).\end{split}$
@@
The proofs of Propositions 10 and 11 are straightforward; we omit the details.
Appendix F The MVPF of Extending Head Start
@@
F.1 The Expected Change in Outcomes
@@ F.2 The MVPF
Appendix G An Example of Positive Codependence
In this section, we provide details for the example that satisfies the positive codependence assumption in Assumption 7. Recall that we have considered the example:
$\mathbb{E}(Y_{i}(2)|u_{i0},u_{i1},u_{i2})-\mathbb{E}Y_{i}(2)=a_{0}u_{i0}+a_{1}u_{i1}+a_{2}u_{i2},$
where $a_{0}$, $a_{1}$, and $a_{2}$ are some constants. Rewrite
$\displaystyle\mathbb{E}(Y_{i}(2)|u_{i0},u_{i1},u_{i2})-\mathbb{E}Y_{i}(2)$ $\displaystyle=(a_{0}+a_{1}+a_{2})u_{i0}-a_{1}(u_{i2}-u_{i1})+(a_{1}+a_{2})(u_{i2}-u_{i0})$
$\displaystyle=(a_{0}+a_{1}+a_{2})u_{i0}-a_{1}\zeta_{i}+(a_{1}+a_{2})\xi_{i}$
$\displaystyle=\mathbb{E}(Y_{i}(2)|u_{i0},\zeta_{i},\xi_{i})-\mathbb{E}Y_{i}(2).$
Hence,
$\displaystyle\mathbb{E}(Y_{i}(2)|\zeta_{i},\xi_{i})-\mathbb{E}Y_{i}(2)$ $\displaystyle=(a_{0}+a_{1}+a_{2})\mathbb{E}(u_{i0}|\zeta_{i},\xi_{i})-a_{1}\zeta_{i}+(a_{1}+a_{2})\xi_{i}.$
Also, recall that we have assumed that $(u_{i0},u_{i1},u_{i2})$ are jointly normal and mutually uncorrelated with the common mean 0 and the common variance 1. Thus,
$\displaystyle\begin{pmatrix}u_{i0}\\ \zeta_{i}\\ \xi_{i}\end{pmatrix}=\begin{pmatrix}1&0&0\\ 0&-1&1\\ -1&0&0\end{pmatrix}\begin{pmatrix}u_{i0}\\ u_{i1}\\ u_{i2}\end{pmatrix}\sim N\left[\begin{pmatrix}0\\ 0\\ 0\end{pmatrix},\begin{pmatrix}1&0&-1\\ 0&2&1\\ -1&1&2\end{pmatrix}\right],$
which implies that
$\displaystyle\mathbb{E}(u_{i0}|\zeta_{i},\xi_{i})$ $\displaystyle=\begin{pmatrix}0&-1\end{pmatrix}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}^{-1}\begin{pmatrix}\zeta_{i}\\ \xi_{i}\end{pmatrix}=\frac{1}{3}\zeta_{i}-\frac{2}{3}\xi_{i}.$
Combining all together yields
$\displaystyle\mathbb{E}(Y_{i}(2)|\zeta_{i},\xi_{i})-\mathbb{E}Y_{i}(2)$ $\displaystyle=(a_{0}+a_{1}+a_{2})\left(\frac{1}{3}\zeta_{i}-\frac{2}{3}\xi_{i}\right)-a_{1}\zeta_{i}+(a_{1}+a_{2})\xi_{i}$
$\displaystyle=\frac{a_{2}+a_{0}-2a_{1}}{3}\zeta_{i}+\frac{a_{2}+a_{1}-2a_{0}}{3}\xi_{i}$
Thus, in this example, Assumption 7 holds if and only if
$a_{2}+a_{0}\geq 2a_{1}\text{ and }a_{2}+a_{1}\geq 2a_{0}.$

### v6

@@ E.1 Notation
$\displaystyle\bm{Q}_{Z}(t)$ $\displaystyle=\left(\mathbb{E}\left(Y\;\mathbf{1}(T=t)\mid Z=z\right)\right)_{z\in\mathcal{Z}}.$
@@ E.2 Theorem T-2 in HP
[…]
(E.2) $\displaystyle\bm{Q}_{S}(t)$ $\displaystyle=\bm{B}_{t}^{\dagger}\bm{Q}_{Z}(t)+\bm{K}_{t}\tilde{\bm{\lambda}}$,
@@
E.3 Identification in the $2\times 3$ model
@@
[…]
@@
[…]
this is not very transparent, however. To derive our identification results, we use (E.2) instead. Note that the equation $\bm{Q}_{S}(t)=\bm{B}_{t}^{\dagger}\bm{Q}_{Z}(t)+\bm{K}_{t}\tilde{\bm{\lambda}}$ holds for any function of $Y(t)$. If we take it to be a constant function of $Y(t)$, we get $\bm{Q}_{S}(t)=\mathbb{E}\bm{S}=\bm{P}_{S}$ and $\bm{Q}_{Z}(t)=\bm{P}_{Z}(t)$, so that (E.2) boils down to
(E.3) $\bm{P}_{S}=\bm{B}_{t}^{\dagger}\bm{P}_{Z}(t)+\bm{K}_{t}\tilde{\bm{\lambda}}\;\text{ for all values of }t.$
@@
Let $(\bm{e}_{s})_{s=1,\ldots,5}$ denote the standard basis vectors in $\mathbb{R}^{5}$. If $\bm{e}^{\prime}_{s}\bm{K}_{t}=0$, then (E.3) point-identifies $\Pr(i\in C^{s})=\bm{e}^{\prime}_{s}\bm{P}_{S}=\bm{e}^{\prime}_{s}\bm{B}_{t}^{\dagger}\bm{P}_{Z}(t)$. Clearly,
@@
this reproduces our identification results for $\bm{P}_{S}$ in Proposition˜5:
@@
Returning to the counterfactual outcomes $Y(t)$, the same argument results in Proposition˜6:
@@
E.4 Identification in the $3\times 3$ model
@@
Let $(\bm{e}_{s})_{s=1,\ldots,8}$ denote the standard basis vectors in $\mathbb{R}^{8}$. Since $\bm{K}_{0}$ has no zero column, none of the $\bm{e}^{\prime}_{s}\bm{K}_{0}$ is zero and the argument in Section E.3 show that no population share $\Pr(i\in C^{s})$ is point-identified by $\bm{B}_{0}^{\dagger}\bm{P}_{Z}(0)$. On the other hand,
@@
and $\bm{e}^{\prime}_{3}\bm{K}_{2}=\bm{e}^{\prime}_{8}\bm{K}_{2}=0$ so that the shares of $C_{222}$ and $C_{212}$ are point-identified. On the other hand, the shares of $C_{010}$, $C_{002}$, and $C_{012}$ are not identified. The results in Propositions 8 and 9 follow.
@@
where we indicate the relevant columns and rows of matrix $\bm{B}_{0}$. Given this pattern, Theorem T-3 and Remark 6.3 in Heckman and Pinto (2018) imply that the unordered monotonicity assumption is not satisfied for the $3\times 3$ model. As mentioned in the main text, this , switching from instrument value from 1 to 2 causes observations in $C_{010}$ to move to treatment $0$, while those in $C_{002}$ move out of treatment $0$. Recall that the ARUM structure rules out “direct two-way flows” (that is, instrument values $1$ and $2$, respectively, make treatments $1$ and $2$ more favorable for everyone). However, the $3\times 3$ model allows for “indirect two-way flows”, where treatment 0 is not targeted by either $z=1$ or $z=2$. Unordered monotonicity is more restrictive than ARUM in that it rules out both direct and indirect two-way flows.
Appendix F The $3\times 3$ Model of Pinto (2022)
@@
(F.1) $\displaystyle\begin{split}\Pr(C_{hhh})&=P(t_{h}|z_{8}),\\ \Pr(C_{lll})&=P(t_{l}|z_{c}),\\ \Pr(C_{mmm})&=P(t_{m}|z_{e}),\\ \Pr(C_{hhm})&=P(t_{h}|z_{e})-P(t_{h}|z_{8}),\\ \Pr(C_{hll})&=P(t_{l}|z_{8})-P(t_{l}|z_{c}),\\ \Pr(C_{mlm})&=P(t_{m}|z_{c})-P(t_{m}|z_{e}),\\ \Pr(C_{hlm})&=1-P(t_{h}|z_{e})-P(t_{l}|z_{8})-P(t_{m}|z_{c}).\end{split}$
@@
(F.2) $\displaystyle\begin{split}P(t_{h}|z_{e})&\geq P(t_{h}|z_{8}),\\ P(t_{l}|z_{8})&\geq P(t_{l}|z_{c}),\\ P(t_{m}|z_{c})&\geq P(t_{m}|z_{e}),\\ 1&\geq P(t_{h}|z_{e})+P(t_{l}|z_{8})+P(t_{m}|z_{c}).\end{split}$
@@
The proofs of Propositions˜13 and 14 are straightforward; we omit the details.
Appendix G The MVPF of Extending Head Start
@@
G.1 The Expected Change in Outcomes
@@ G.2 The MVPF
Appendix H A Roy Model with Positive Selection
In this section, we provide details for the example in Section 3.3.3. Recall that we assumed that
$\mathbb{E}(Y_{i}(2)\mid u_{i0},u_{i1},u_{i2})-\mathbb{E}(Y_{i}(2))=a_{0}u_{i0}+a_{1}u_{i1}+a_{2}u_{i2}\equiv v_{i},$
where $a_{0}$, $a_{1}$, and $a_{2}$ are some constants. Since $(u_{i0},u_{i1},u_{i2})$ are jointly normal, taking expectations conditional on $u_{i2}-u_{i1}=\zeta_{i}$ and $u_{i2}-u_{i0}=\xi_{i}$ gives
$\mathbb{E}(Y_{i}(2)\mid\zeta_{i},\xi_{i})-\mathbb{E}(Y_{i}(2))=\begin{pmatrix}\text{Cov}(v_{i},\zeta_{i})&\text{Cov}(v_{i},\xi_{i})\end{pmatrix}\left(\mathbb{V}\mathrm{ar}\begin{pmatrix}\zeta_{i}\\ \xi_{i}\end{pmatrix}\right)^{-1}\begin{pmatrix}\zeta_{i}\\ \xi_{i}\end{pmatrix}$.
Simple calculations give
$\mathbb{E}(Y_{i}(2)\mid\zeta_{i},\xi_{i})-\mathbb{E}(Y_{i}(2))=\frac{a_{2}+a_{0}-2a_{1}}{3}\zeta_{i}+\frac{a_{2}+a_{1}-2a_{0}}{3}\xi_{i}.$
The coefficients of $\zeta_{i}$ and $\xi_{i}$ are non-negative if and only if
[…]
which can be written more compactly as $a_{2}\geq\max(a_{0},a_{1})+|a_{1}-a_{0}|$. Assume that this inequality holds, and consider Figure 5. Every point in $C_{112}^{(i)}$ has lower values of both $\zeta_{i}$ and $\xi_{i}$ than any point in $C_{212}$. Therefore, the expected value of $Y_{i}(2)$ in this triangle is smaller than $\mathbb{E}(Y_{i}(2)\mid i\in C_{212})$. Every point in $C_{112}^{(ii)}$ has a smaller value of $\zeta_{i}$ than any point in $C_{212}$ (fixing the value of $\xi_{i}\geq-\underline{U}(2)$ on both sides). The expected value of $Y_{i}(2)$ in this rectangle is again smaller than $\mathbb{E}(Y_{i}(2)\mid i\in C_{212})$. Combining these two inequalities gives $\mathbb{E}(Y_{i}(2)\mid i\in C_{212})\geq\mathbb{E}(Y_{i}(2)\mid i\in C_{112})$, that is (3.10).